What’s the 50th term of this linear sequence 27,25,23,21,19...
1
Expert's answer
2019-09-23T09:35:59-0400

The linear sequence 27, 25, 23, 21, 19... is an arithmetic progression with (-2) as the common difference.

The nth term of an arithmetic progression is "a_n=a_1+(n-1)d", where "a_n" is the nth term, "a_1" is the first term, d is the common difference.

In our case, "n=50, a_1=27, d=-2" , hence, we get

"a_{50}=27+49\cdot(-2);\\
a_{50}= 27-98=-71."

The 50th term of this sequence is (-71).